# On the Estimation of bivariate Conditional Transition Rates

### v1

@@ Abstract
Recent literature has found conditional transition rates to be a useful tool for avoiding Markov assumptions in multistate models. While the estimation of univariate conditional transition rates has been extensively studied, the intertemporal dependencies captured in the bivariate conditional transition rates still require a consistent estimator. We provide an estimator that is suitable for censored data and emphasize the connection to the rich theory of the estimation of bivariate survival functions. Bivariate conditional transition rates are necessary for various applications in the survival context but especially in the calculation of moments in life insurance mathematics.
Keywords: bivariate survivor function; censoring; non-Markov models; non-parametric estimation; life & health insurance; free-policy option
@@ 1 Introduction
In survival statistics, the Kaplan-Meier estimator, as seen in Kaplan & Meier [18], is widely used to obtain non-parametric estimation of survival functions for various modelling purposes when censoring is involved. This applies to various scenarios, such as life insurance modelling, time to machine failure, or the time to illness after the application of medicine. However, this estimator can only take into account a single event and therefore cannot capture dependencies between different events included in the data. Consequently, its suitability for analysing data sets, particularly those arising from genetic epidemiological studies, for disease events of family members, or other similar dependencies, is limited. Many different estimators have been proposed to fill this gap. Dabrowska [10] uses two-dimensional product integrals to estimate the bivariate survival function, whilst Prentice & Cai [20] employ Volterra integral equation techniques to connect the covariance function and the marginal survivor functions with the bivariate survivor function. Van der Laan [23] uses non-parametric maximum likelihood estimation, and recently, Prentice & Zhao [21] extended the work of Dabrowska with a recursive formula for arbitrary numbers of failure times. Gill [13] gives a reason why there are so many different estimators: the fact that in two and higher dimensions there are different ways to get from one time point to another time point.
@@
Another important concept in survival analysis is the notion of recurrent events used to study recurring episodes of infection or hospitalisation. This is often modelled by introducing multistate models, described by Hougaard [17], which are mostly based on modelling with Markov assumptions. In these models, estimators often target cumulative intensity processes which provide insight into the expected number of upcoming transitions. These estimators are now known as the Nelson-Aalen estimator for transition rates and the Aalen-Johansen estimator of transition probabilities, see Aalen [1] and Borgan, Gill & Keiding [6]. Datta & Satten [11] discovered that the Nelson-Aalen and Aalen-Johansen estimators remain consistent on data that does not follow the Markov assumption. However, cumulative transition rates and transition probabilities do not hold full information about these state processes without the Markov assumption, as they do not capture the temporal interdependencies between consecutive jump events, necessary, for example, for calculating variances or higher moments. To fill this gap, Bathke & Christiansen [4] have recently proposed a novel framework involving bivariate transition rates but without suitable non-parametric estimators. This paper aims to address the lack of estimation procedures for general bivariate intensity rates. We will demonstrate the relationship to the estimation of bivariate survival functions, particularly the works of Dabrowska [10] and Prentice & Cai [20]. Our aim is to construct an estimator and demonstrate its consistency through almost sure uniform convergence. This is a widely accepted method for understanding the asymptotic properties of estimators, as demonstrated by Shorack & Wellner [22].
@@
Bivariate rates have another important application apart from being used for classical survival analysis applications. They are also key for the calculation of life insurance premiums or reserves of insurance contracts with special payment functions in non-Markov models. Insurance cash flows typically include transition payments from one state to another, as well as sojourn payments for remaining in the corresponding state. Transition payments can be modelled using jump processes, while sojourn payments can be modelled using indicator functions. Guibert & Planchet [16] were able to estimate reserves in a general non-Markov multi-state model for sojourn payments. Christiansen [7] extended this to reserves of insurance cash flows with both transition and sojourn payments, using path-independent payment functions. He showed that the consistency properties of the estimators can be transferred to reserve estimation. Our contribution is to show that our bivariate estimators of transition rates and probabilities can be used to estimate reserves for specific path-dependent cash flows and second moments of future liabilities, which are important for many types of risk analysis.
@@
The paper is structured as follows. In Section 2 the general multi-state modelling framework is introduced, here we connect jump processes and indicator processes. In Section 3 we will introduce our bivariate notation. Section 4 introduces the definition of bivariate conditional rates and the corresponding integral equations, which will be connected to bivariate Peano-series. Section 5 focuses on estimating the bivariate conditional rates using a bivariate landmark Nelson-Aalen estimator, as well as the bivariate transition probabilities using the bivariate landmark Aalen-Johansen estimator. Section 6 demonstrates how these estimators can be utilised to form estimators for cumulative expected values of general cash flows with interdependencies between two time points in the context of life insurance.
@@ 2 Multi-state modelling framework
which describes the current state that the individual is in. This process is modelled as an adapted càdlàg jump process on a finite state space $\mathcal{Z}$. Additionally, we set $Z$ to be continuous at 00, i.e. $Z_{0-}:=Z_{0}$. When modelling these objects, we want to be able to describe past, presence, and future developments. Thus, we assume that we are currently at time $s\geq 0$. So the time interval $[0,s]$ represents the past and the present, and the time interval $(s,\infty)$ represents the future. This time point $s$ is arbitrary but fixed throughout this paper. Based on $Z$ we define additional processes, first the state indicator processes $(I_{i})_{i\in\mathcal{Z}}$ by
@@
Currently, Markov modelling is widely used in practice. Therefore, most of the available data is of Markov type. Thus, the natural extension would be the as-if Markov model, as it combines classical Markov information with a non-Markov model. In addition, the numerical complexity will be increased whenever we work with a broader information model. Thus, we will focus on an estimation procedure that works with discrete covariates, which includes the as-if Markov model because #$\#\mathcal{Z}<\infty$.
@@ 3 Notation
For the two-dimensional variables, definitions of processes and integrals, we use $\boldsymbol{t}=(t_{1},t_{2})$, $f(\boldsymbol{u}^{-})=f(u_{1}^{-},u_{2}^{-})$ for a bivariate upper continuous function $f$ that has uniform lower limits, and $f(\mathrm{d}\boldsymbol{u})=f(\mathrm{d}u_{1},\mathrm{d}u_{2})$ for a two-variable function f with finite two-dimensional total variation, see the Appendix A.5 for the definition of the total variation used in this paper. In addition, we use partial ordering on $\mathbb{R}^{2}$, meaning that $\boldsymbol{x}\leq\boldsymbol{y}$ if and only if $x_{1}\leq y_{1}$ and $x_{2}\leq y_{2}$. We extend this notation so that $\boldsymbol{x}$ can be set into relation with a one-dimensional constant. For example, $\boldsymbol{x}\leq y$ iff $x_{1}\leq y$ and $x_{2}\leq y$. Consequently, $(s,\boldsymbol{t}]$ is a two-dimensional rectangle that is open at the bottom and the left boundaries and closed at the top and right boundaries with the corner points $(s,s)$ and $(t_{1},t_{2})$. Apart from two-dimensional variables, we will use a lot of indices to model different jumps at different times. For this we introduce the two-dimensional indices $\boldsymbol{i}=(i_{1},i_{2})$.
4 Conditional Transition Rate
@@
Let $P_{z,i}=(P_{z,i}(t))_{t\geq 0}$ and $Q_{z,ij}=(Q_{z,ij}(t))_{t\geq 0}$ for $i,j\in\mathcal{Z}$ be the almost surely unique càdlàg paths that satisfy
@@
Let $\boldsymbol{P}_{z,\boldsymbol{i}}=(\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{t}))_{t_{1},t_{2}\geq 0}$ and $\boldsymbol{Q}_{z,\boldsymbol{i}\boldsymbol{j}}=(\boldsymbol{Q}_{z,\boldsymbol{i}\boldsymbol{j}}(\boldsymbol{t}))_{t_{1},t_{2}\geq 0}$ for $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$ be the almost surely unique surfaces that are càdlàg in each variable and satisfy
@@
For $i,j\in\mathcal{Z}$ and $t\in(s,\infty)$ let the stochastic process $(\Lambda_{ij}(t))_{t\geq s}$ be defined by
@@
and for $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$ and $\boldsymbol{t}\in(s,\infty)^{2}$ let the random surface $(\boldsymbol{\Lambda}_{\boldsymbol{i}\boldsymbol{j}}(\boldsymbol{t}))_{\boldsymbol{t}\geq s}$ be defined by
@@
$\begin{split}\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{t})&=\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{s})+I_{i_{2}}(s)\left(P_{z,i_{1}}(t_{1})-P_{z,i_{1}}(s)\right)+I_{i_{1}}(s)\left(P_{z,i_{2}}(t_{2})-P_{z,i_{2}}(s)\right)\\ &\quad+\;\int\displaylimits_{(s,\boldsymbol{t}]}\sum_{\begin{subarray}{c}\boldsymbol{j}\in\mathcal{Z}^{2}\end{subarray}}{\boldsymbol{P}}_{z,\boldsymbol{j}}(\boldsymbol{u}^{-}){\boldsymbol{\Lambda}}_{z,\boldsymbol{j}\boldsymbol{i}}(\mathrm{d}\boldsymbol{u}),\end{split}$ (4.5)
for $\boldsymbol{t}\geq s$ and $\boldsymbol{i}=(i_{1},i_{2})\in\mathcal{Z}^{2}$. These integral equations are similar to the inhomogeneous Volterra equations found in the survival setting, see Prentice and Cai [20].
@@ 5 Bivariate Conditional Nelson-Aalen and Aalen-Johansen Estimation
We begin this section with the general estimation setup. For the rest of this paper we will assume that we are observing $n\in\mathbb{N}$ individuals:
@@
    $R^{1}$ is stochastically independent of $(Z^{1}(t))_{t\geq s}$ and almost surely greater then $s$.
This means that we assume that $\xi^{m}$ is not affected by censoring. That is, it is observable in the data at time $s$ for all individuals.
@@
for $j,k\in\mathcal{Z}$ and $t>s$. This in turn allow us to define censored conditional transition rates as
@@
and the corresponding bivariate censored conditional transition rates:
@@
In the rest of this paper, we will always assume that $\boldsymbol{\tau}$ satisfies this condition and is greater than $s$. We have difficulties estimating a process like $\boldsymbol{\Lambda}^{c}_{z,\boldsymbol{ij}}(\boldsymbol{t})$ in the case that $\boldsymbol{P}^{c}_{z,\boldsymbol{i}}(\boldsymbol{u}^{-})$ is close to 00. There are different strategies to circumvent this problem. See for example IV.1.2 from Andersen, Borgan and Gill [3] where they use assumptions on the behaviour of jump processes to circumvent this problem. For the estimation of bivariate survival functions this gets solved by the fact that $P^{c}(u-)$ goes to zero monotonously as $u$ goes to infinity, which does not generally happen in a multi-state model. We will use the perturbation idea of Bladt & Furrer [5], which we adapt from the univariate case. This corresponds to the new process $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}(\boldsymbol{t})$ with
@@
for $\epsilon>0$, $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$ and $\boldsymbol{t}>s$. This $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}(\boldsymbol{t})$ will be estimated as a substitute to $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\boldsymbol{t})$.
@@
which means that we will get arbitrarily close to the unperturbed conditional transition rates. Second, for practical calculations and estimates, we generally use discretisation of our conditional transition rates and probabilities, i.e. a discretisation in the time and value axis. This means that the only way to get infinitely close to zero is to make a jump to zero. These jumps to zero are not a problem if $\epsilon$ is taken to be smaller than the smallest jump. See Bladt & Furrer [5] example after Definition 2.4 for similar reasoning in the case of perturbation of the univariate conditional transition rates. Furthermore, in practice a complete convergence of transition probabilities to zero can be stopped anyway by a forced transition of states. Examples of this are transitions to annuity at certain ages. In implementation, this leads to forced transitions, for example, into the state ”dead” or ”retired”. These forced transitions prevent the different transition probabilities from converging to zero. So in practice we usually get
@@
for $t\in(s,\boldsymbol{\tau}]$, $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$, and a fixed $\epsilon>0$.
Definition 5.4.
We start by defining the estimators for the expected value of censored jump processes and indicator processes.
@@
and bivariate landmark Aalen-Johansen estimators as $P_{z,ik}^{(n,\epsilon)}$, the unique solution to the integral equation
$\begin{split}\boldsymbol{P}_{z,\boldsymbol{i}}^{(n,\epsilon)}(\boldsymbol{t})&=\boldsymbol{I}_{z,\boldsymbol{i}}^{(n)}(s)+I_{z,i_{2}}^{(n)}(s)\left(P_{z,i_{1}}^{(n,\epsilon)}(t_{1})-I_{z,i_{1}}^{(n)}(s)\right)+I_{z,i_{1}}^{(n)}(s)\left(P_{z,i_{2}}^{(n,\epsilon)}(t_{2})-I_{z,i_{2}}^{(n)}(s)\right)\\ &\quad+\int\displaylimits_{(s,\boldsymbol{t}]}\sum_{\begin{subarray}{c}\boldsymbol{j}\in\mathcal{Z}^{2}\end{subarray}}{\boldsymbol{P}}^{(n,\epsilon)}_{z,\boldsymbol{j}}(\boldsymbol{u}^{-}){\boldsymbol{\Lambda}}^{(n,\epsilon)}_{z,\boldsymbol{ji}}(\mathrm{d}\boldsymbol{u}),\end{split}$ (5.1)
@@
The proof for this theorem will follow at the end of this section.
@@
The proof for this lemma will follow at the end of this section.
@@
The proof for this theorem will follow at the end of this section.
This theorem somewhat describes consistency of the estimators $\boldsymbol{\Lambda}^{(n,\epsilon)}_{z,\boldsymbol{ij}}$ and $\boldsymbol{P}_{z}^{(n,\epsilon)}$. In practice one can fix an $\epsilon$ and let $n$ grow to get a reasonably close estimation of $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}$ and $\boldsymbol{P}_{z}$.

### v2

@@ Abstract
Recent literature has found conditional transition rates to be a useful tool for avoiding Markov assumptions in multi-state models. While the estimation of univariate conditional transition rates has been extensively studied, the intertemporal dependencies captured in the bivariate conditional transition rates still require a consistent estimator. We provide an estimator that is suitable for censored data and emphasize the connection to the rich theory of the estimation of bivariate survival functions. Bivariate conditional transition rates are necessary for various applications in the survival context but especially in the calculation of moments in life insurance mathematics.
Keywords: bivariate survivor function; censoring; non-Markov models; non-parametric estimation; life & health insurance; free policy option
@@ 1 Introduction
In survival statistics, the Kaplan-Meier estimator, as seen in Kaplan & Meier [21], is widely used to obtain non-parametric estimation of survival functions for various modelling purposes when censoring is involved. This applies to various scenarios, such as life insurance modelling, time to machine failure, or the time to illness after the application of medicine. However, this estimator can only take into account a single event and therefore cannot capture dependencies between different events included in the data. Consequently, its suitability for analysing data sets, particularly those arising from genetic epidemiological studies, for disease events of family members, or other similar dependencies, is limited. Many different estimators have been proposed to fill this gap. Dabrowska [13] uses two-dimensional product integrals to estimate the bivariate survival function, whilst Prentice & Cai [24] employ Volterra integral equation techniques to connect the covariance function and the marginal survivor functions with the bivariate survivor function. Van der Laan [30] uses non-parametric maximum likelihood estimation, and recently, Prentice & Zhao [25] extended the work of Dabrowska with a recursive formula for arbitrary numbers of failure times. Gill [16] gives a reason why there are so many different estimators: the fact that in two and higher dimensions there are different ways to get from one point in the two-dimensional plane to another point in the two-dimensional plane.
@@
Another important concept in survival analysis is the notion of recurrent events used to study recurring episodes of infection or hospitalisation. This is often modelled by introducing multi-state models, described by Hougaard [20], which are mostly based on modelling with Markov assumptions. In these models, estimators often target cumulative intensity processes which provide insight into the expected number of upcoming transitions. These estimators are now known as the Nelson-Aalen estimator for transition rates and the Aalen-Johansen estimator of transition probabilities, see Aalen [1] and Borgan, Gill & Keiding [8]. Datta & Satten [14] discovered that the Nelson-Aalen and Aalen-Johansen estimators remain consistent on data that does not follow the Markov assumption. The use case for multi-state models without the Markov assumption, referred to here as non-Markov models, has been showcased by Putter and Spitoni in [26]. Their simulation studies demonstrate that, in some instances, Markov estimation introduces a bias when the data does not follow the Markov assumption. In this case, there is a favourable bias-variance trade-off when using the landmark Nelson-Aalen and Aalen-Johansen estimators, which target transition rates and probabilities of a non-Markov model.
@@
However, transition rates and transition probabilities do not hold full information about the distribution of these state processes without the Markov assumption, as they do not capture the temporal interdependencies between consecutive jump events, necessary, for example, for calculating variances or higher moments. To fill this gap, Bathke & Christiansen [5] have recently proposed a novel framework involving bivariate transition rates but without suitable non-parametric estimators. This paper aims to address the lack of estimation procedures for general bivariate intensity rates. We demonstrate the relationship to the estimation of bivariate survival functions, particularly the works of Dabrowska [13] and Prentice & Cai [24]. Our aim is to construct a landmark estimator and demonstrate its consistency through almost sure uniform convergence. This is a widely accepted method for understanding the asymptotic properties of estimators, as demonstrated by Shorack & Wellner [28]. Bivariate rates have another important application apart from being used for classical survival analysis applications. They are also key for the calculation of life insurance premiums or reserves of insurance contracts with special payment functions in non-Markov models. This is important because life insurance calculation is generally done in Markov or semi-Markov models, which introduce systematic model risk. Insurance cash flows typically include transition payments from one state to another, as well as sojourn payments for remaining in the corresponding state. Transition payments can be modelled using jump processes, while sojourn payments can be modelled using indicator functions. Guibert & Planchet [19] were able to estimate reserves in a general non-Markov multi-state model for sojourn payments. Christiansen [10] extended this to reserves of insurance cash flows with both transition and sojourn payments, using path-independent payment functions. He showed that the consistency properties of the estimators can be transferred to reserve estimation. Our contribution is to show that our bivariate estimators of transition rates and probabilities can be used to estimate reserves for specific path-dependent cash flows and second moments of future liabilities, which are important for many types of risk analysis.
@@
The paper is structured as follows. In Section 2 the general multi-state modelling framework is introduced, here we connect jump processes and indicator processes. In Section 3 we introduce our bivariate notation. Section 4 introduces the definition of bivariate conditional rates and the corresponding integral equations, which will be connected to bivariate Peano-series. Section 5 focuses on estimating the bivariate conditional rates using a bivariate landmark Nelson-Aalen estimator, as well as the bivariate transition probabilities using the bivariate landmark Aalen-Johansen estimator. In this Section we introduce Theorem 5.9 which is the main contribution of this paper. Section 6 demonstrates how these estimators can be utilised to form estimators for cumulative expected values of general cash flows with interdependencies between two time points in the context of life insurance. Section 7 introduces a numerical example, where the set up we introduce in Section 6 is used to form estimators for expected future cash flows in the case of scaled payments in life insurance.
@@ 2 Multi-state modelling framework
which describes the current state that the individual is in. This process is modelled as an adapted càdlàg jump process on a finite state space $\mathcal{Z}$. When modelling these objects, we want to be able to describe past, present, and future developments. Thus, we assume that we are currently at time $s\geq 0$. So the time interval $[0,s]$ represents the past and the present, and the time interval $(s,\infty)$ represents the future. This time point $s$ is arbitrary but fixed throughout this paper. Based on $Z$ we define additional processes, first the state indicator processes $(I_{i})_{i\in\mathcal{Z}}$ by
@@
Currently, Markov modelling is widely used in practice. Therefore, most of the available data is of Markov type. Thus, the natural extension would be the as-if Markov model, as it combines classical Markov information with a non-Markov model. In addition, the numerical complexity increases whenever we work with a broader information model. We therefore focus on an estimation procedure that conditions on discrete random variables, which includes the as-if Markov model because #$\#\mathcal{Z}<\infty$.
@@ 3 Notation
For two-dimensional variables, definitions of processes and integrals, we use $\boldsymbol{t}=(t_{1},t_{2})$, $f(\boldsymbol{u}^{-})=f(u_{1}^{-},u_{2}^{-})$ for a bivariate upper continuous function $f$ that has uniform lower limits, and $f(\mathrm{d}\boldsymbol{u})=f(\mathrm{d}u_{1},\mathrm{d}u_{2})$ for a two-variable function $f$ with finite two-dimensional total variation, see the Appendix A.5 for the definition of the total variation used in this paper. In addition, we use partial ordering on $\mathbb{R}^{2}$, meaning that $\boldsymbol{x}\leq\boldsymbol{y}$ if and only if $x_{1}\leq y_{1}$ and $x_{2}\leq y_{2}$. We extend this notation so that $\boldsymbol{x}$ can be set into relation with a one-dimensional constant. For example, $\boldsymbol{x}\leq y$ iff $x_{1}\leq y$ and $x_{2}\leq y$. Consequently, $(s,\boldsymbol{t}]$ is a two-dimensional rectangle that is open at the bottom and the left boundaries and closed at the top and right boundaries with the corner points $(s,s)$ and $(t_{1},t_{2})$. Apart from two-dimensional variables, we use various indices to model different jumps at different times. For this we introduce the two-dimensional indices $\boldsymbol{i}=(i_{1},i_{2})$.
4 Conditional Transition Rates
@@
Let $P_{z,i}=(P_{z,i}(t))_{t>s}$ and $Q_{z,ij}=(Q_{z,ij}(t))_{t>s}$ for $i,j\in\mathcal{Z}$ be the almost surely unique càdlàg paths that satisfy
@@
Let $\boldsymbol{P}_{z,\boldsymbol{i}}=(\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{t}))_{t_{1},t_{2}>s}$ and $\boldsymbol{Q}_{z,\boldsymbol{i}\boldsymbol{j}}=(\boldsymbol{Q}_{z,\boldsymbol{i}\boldsymbol{j}}(\boldsymbol{t}))_{t_{1},t_{2}>s}$ for $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$ be the almost surely unique surfaces that are càdlàg in each variable and satisfy
@@
Remark 4.1.
Following section 2, $\xi=Z(s)$ is a natural example of a landmark, as seen for example in the as-if Markov model. In this case, $z\in\mathcal{Z}$. But the landmark can also include discrete external landmarks such as gender. So, in general, $z$ need not to be in $\mathcal{Z}$. The fact that the landmark $\xi$ is discrete is important for the landmarking ideas we use in this paper. See [7] for an approach to conditioning on continuous information.
@@
For $i,j\in\mathcal{Z}$ and $t\in(s,\infty)$ let the function $(\Lambda_{z,ij}(t))_{t\geq s}$ be defined by
@@
and for $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$ and $\boldsymbol{t}\in(s,\infty)^{2}$ let the function $(\boldsymbol{\Lambda}_{z,\boldsymbol{i}\boldsymbol{j}}(\boldsymbol{t}))_{\boldsymbol{t}\geq s}$ be defined by
@@
$\begin{split}\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{t})&=\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{s})+P_{z,i_{2}}(s)\left(P_{z,i_{1}}(t_{1})-P_{z,i_{1}}(s)\right)+P_{z,i_{1}}(s)\left(P_{z,i_{2}}(t_{2})-P_{z,i_{2}}(s)\right)\\ &\quad+\;\int\displaylimits_{(s,\boldsymbol{t}]}\sum_{\begin{subarray}{c}\boldsymbol{j}\in\mathcal{Z}^{2}\end{subarray}}{\boldsymbol{P}}_{z,\boldsymbol{j}}(\boldsymbol{u}^{-}){\boldsymbol{\Lambda}}_{z,\boldsymbol{j}\boldsymbol{i}}(\mathrm{d}\boldsymbol{u}),\end{split}$ (4.5)
for $\boldsymbol{t}\geq s$ and $\boldsymbol{i}=(i_{1},i_{2})\in\mathcal{Z}^{2}$.
These integral equations are similar to the inhomogeneous Volterra equations found in the survival setting, see Prentice and Cai [24].
@@ 5 Bivariate Conditional Nelson-Aalen and Aalen-Johansen Estimation
We begin this section with the general non-parametric estimation setup, which is similar to the non-Markov setup in Christiansen [10]. The non-parametric approach of this paper is deeply rooted in the fact that many practitioners, for example in the German actuarial community, use purely non-parametric estimators. Moreover, the non-parametric approach can be used as a preview to better calibrate parametric models. First, we present the general estimation setup.
For the rest of this paper we assume that we are observing $n\in\mathbb{N}$ individuals:
@@
    $R^{1}$ is stochastically independent of $(Z^{1}(t))_{t\geq s}$ and $\xi^{1}$ and almost surely greater than $s$.
It means that we assume that $\xi^{m}$ is not affected by censoring. That is, it is observable in the data at time $s$ for all individuals.
Remark 5.2.
The estimation model could be defined exactly as Christiansen does in [10], but we think that left censoring and weaker conditions on right censoring are not worth the notational overhead. One main application of censoring or truncation in the actuarial context in section 6 is that the portfolio of insured, and thus the data we use for estimation, is not fully closed, so we have data that may be right censored but is rarely left censored.
Remark 5.3.
There are many other approaches to the modelling of censoring than the one we present in this paper. There is the example of inverse probability of censoring weighting methods, which allow for additional dependency between the jump process and the censoring mechanism. This in turn requires a parametric censoring model, which distinguishes their approach from our fully non-parametric one, as already mentioned by Bladt and Furrer in [7]. In addition, current work on IPCW methods in multi-state models uses a simpler setup, see the work of Mostajabi and Datta [23] and Siriwardhana, Kulasekera and Datta [29].
@@
for $j,k\in\mathcal{Z}$ and $t>s$. This in turn allows us to define censored transition rates as
@@
and the corresponding bivariate conditional transition rates:
@@
In the rest of this paper, we always assume that $\boldsymbol{\tau}$ satisfies this condition and is greater than $s$. In practice, $\tau$ can be understood as the right endpoint of the data window, we are able to observe and is mostly predetermined by the application. We have difficulties estimating a process like $\boldsymbol{\Lambda}^{c}_{z,\boldsymbol{ij}}(\boldsymbol{t})$ in the case that $\boldsymbol{P}^{c}_{z,\boldsymbol{i}}(\boldsymbol{u}^{-})$ is close to 00. There are different strategies to circumvent this problem. See for example IV.1.2 from Andersen, Borgan and Gill [4] where they use assumptions on the behaviour of jump processes to circumvent this problem. For the estimation of bivariate survival functions this gets solved by the fact that $P^{c}(u-)$ goes to zero monotonically as $u$ goes to infinity, which does not generally happen in a multi-state model. We use the perturbation idea of Bladt & Furrer [7], which we adapt from the univariate case. This corresponds to the new process $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}(\boldsymbol{t})$ with
@@
for $\epsilon>0$, $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$ and $\boldsymbol{t}>s$. This $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}(\boldsymbol{t})$ is estimated as a replacement for $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\boldsymbol{t})$.
@@
which means that we get arbitrarily close to the unperturbed conditional transition rates. Second, for practical calculations and estimates, we generally use discretisation of our conditional transition rates and probabilities, i.e. a discretisation in the time and value axis. This means that the only way to get infinitely close to zero is to make a jump to zero. These jumps to zero are not a problem if $\epsilon$ is taken to be smaller than the smallest jump. See Bladt & Furrer [7] example after Definition 2.4 for similar reasoning in the case of perturbation of the univariate conditional transition rates. Furthermore, in practice a complete convergence of transition probabilities to zero can be stopped anyway by a forced transition of states. Examples of this are transitions to annuity at certain ages. In implementation, this leads to forced transitions, for example, into the state ”dead” or ”retired”. These forced transitions prevent the different transition probabilities from converging to zero. So in practice we usually get
@@
for $t\in(s,\boldsymbol{\tau}]$, $\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}^{2}$, and a fixed $\epsilon>0$. We start by defining the estimators for the expected value of censored jump processes and indicator processes.
@@
and bivariate landmark Aalen Johansen estimators as $P_{z,ik}^{(n,\epsilon)}$, the unique solution to the integral equation
$\begin{split}\boldsymbol{P}_{z,\boldsymbol{i}}^{(n,\epsilon)}(\boldsymbol{t})&=\boldsymbol{P}_{z,\boldsymbol{i}}(s)+P_{z,i_{2}}(s)\left(P_{z,i_{1}}^{(n,\epsilon)}(t_{1})-P_{z,i_{1}}(s)\right)+P_{z,i_{1}}(s)\left(P_{z,i_{2}}^{(n,\epsilon)}(t_{2})-P_{z,i_{2}}(s)\right)\\ &\quad+\int\displaylimits_{(s,\boldsymbol{t}]}\sum_{\begin{subarray}{c}\boldsymbol{j}\in\mathcal{Z}^{2}\end{subarray}}{\boldsymbol{P}}^{(n,\epsilon)}_{z,\boldsymbol{j}}(\boldsymbol{u}^{-}){\boldsymbol{\Lambda}}^{(n,\epsilon)}_{z,\boldsymbol{ji}}(\mathrm{d}\boldsymbol{u}),\end{split}$ (5.1)
@@
Remark 5.6.
In the construction of the estimation model, we assume that we observe individuals until they are censored. The construction of the landmark Nelson-Aalen estimator does not depend on this knowledge. We only need to observe the individuals until they reach an absorption time, i.e. they jump into an absorption state. This is the case because there can be no jumps after the absorption time, so $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(n,\epsilon)}(\boldsymbol{t})$ does not depend on this information. And since the bivariate landmark Aalen-Johansen estimator only depends on the landmark Nelson-Aalen estimator, this information is not really needed.
@@
The proof for this theorem follows at the end of this section.
@@
The proof for this lemma follows at the end of this section.
@@
The proof for this theorem follows at the end of this section.
This theorem somewhat describes consistency of the estimators $\boldsymbol{\Lambda}^{(n,\epsilon)}_{z,\boldsymbol{ij}}$ and $\boldsymbol{P}_{z}^{(n,\epsilon)}$. In practice one can fix an $\epsilon$ and let $n$ grow to get a reasonably close estimation of $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}$ and $\boldsymbol{P}_{z}$. In the proof of this theorem we show, among other things, that the solution of the equation (A.1) as a functional of $\boldsymbol{\Lambda}_{z,\boldsymbol{i}\boldsymbol{j}}$ and $\Lambda_{z,ij}$ is Hadamard continuous. This statement is already known, see Gill, van der Laan & Wellner [17]. They did not add a complete proof, so we include this part in our proof. Additionally, they state that this functional is Hadamard differentiable.